OpenStudy (anonymous):

HI can anyone help me? :D :D <3 I am a little worried that i might have this question wrong . if you help i will totally give you a medal and become your fan !! 1:2 = 4:8 The product of both the means and the extremes is my answer is 2 because 2 goes into 4 and 8, am I correct ?

OpenStudy (anonymous):

they are in proportion 8:8ur right

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

why did you put 8:8

OpenStudy (anonymous):

@ChiefArnav

OpenStudy (anonymous):

cuz 1:2::4:8 means (1 x 8) and (2 x 4) which is 8:8

OpenStudy (anonymous):

r u asked to simplify?

OpenStudy (anonymous):

oh ! so 2 is wrong because my answer choices are a.2 b.4 c.8 d.15 @ChiefArnav

OpenStudy (anonymous):

r u asked to simplify?

OpenStudy (anonymous):

go with c

OpenStudy (anonymous):

okay thanks

OpenStudy (anonymous):

your welcome

OpenStudy (anonymous):

more question?

OpenStudy (anonymous):

i have one more question that i really dont understand so well

OpenStudy (anonymous):

ask

OpenStudy (anonymous):

All of the following proportions are equivalent except a/b=c/d b/a=d/c a/d=c/b a/c=b/d

OpenStudy (anonymous):

alright thank you for all your helpand time!

OpenStudy (anonymous):

happy was able to help :3

OpenStudy (anonymous):

im so srry its 2 and 3

OpenStudy (anonymous):

not 1 and 4